OpenStudy (anonymous):

which expression is a polynomial? a. 11/√x b. √11x c. x - √11 d. 11/x^2 I need help asap!

OpenStudy (anonymous):

In a polynomial : x cannot be under Square Root or Denominator, power of x should be positive..

OpenStudy (anonymous):

So its C..?

OpenStudy (anonymous):

Good...

OpenStudy (anonymous):

thank you can you please help me with one more?

OpenStudy (anonymous):

Simplify the expression. 34 • 32 (Points : 4) 36 38 96 98

OpenStudy (anonymous):

3^4 * 3^2

OpenStudy (anonymous):

Do you know the meaning of ^ ??

OpenStudy (anonymous):

Yeah the answers are 3^6, 3^8, 9^6, 9^8 sorry my computer didnt show it

OpenStudy (anonymous):

Yeah I know how it works i'm just confused on this question

OpenStudy (anonymous):

So, your question is this: \((3)^4. (3)^2\) = ??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

See, if number is same, then we have these two identities: \[\large (a)^b \cdot (a)^c = (a)^{b+c}\] \[\large \frac{(a)^b}{(a)^c} = (a)^{b-c}\]

OpenStudy (anonymous):

Use the first one to get your answer..

OpenStudy (anonymous):

729..?

OpenStudy (anonymous):

Use the first one.. Don't solve it.. you have a = 3, b = 4 and c = 2, just put it in first identity..

OpenStudy (anonymous):

See, you have to just add the powers, that's it..

OpenStudy (anonymous):

a. 3^6 ?

OpenStudy (anonymous):

Going good..

OpenStudy (anonymous):

\(\huge \color{green}{\mathsf{Welcome \; \; To \; \; Openstudy...}}\)

OpenStudy (anonymous):

I get it now thank you!!